OpenStudy (anonymous):

Am I right? If r(x) = -4x - 1 and t(x) = 2x - 1, find (r + t)(7). Is -16?

OpenStudy (hoblos):

yes you are :D

OpenStudy (anonymous):

Thank god!! James really did help me a lot!

OpenStudy (hoblos):

:D:D

OpenStudy (anonymous):

you see, you can do it yourself

OpenStudy (anonymous):

If h(x) = x - 1 and j(x) = -4x, find h[j(5)]. Is -16 as well?

OpenStudy (hoblos):

no this is -21

OpenStudy (anonymous):

o-o;; But.. What did I do wrong then?

OpenStudy (jamesj):

good, for one problem. Now the other. ======== If h(x) = x - 1 and j(x) = -4x, find h[j(5)]. Find first the value of j(5)

OpenStudy (jamesj):

j(5) = .... what?

OpenStudy (anonymous):

-20

OpenStudy (jamesj):

Right. Hence h(j(5)) = h(20) = ... what?

OpenStudy (anonymous):

... You lost me there.

OpenStudy (jamesj):

You want to find the value of h(j(5)). Now we know that j(5) = -20. Therefore h( j(5) ) = h( -20 ) as we just substituted the value for j(5)

OpenStudy (anonymous):

5(-20)= -100

OpenStudy (jamesj):

No. The function h is defined as h(x) = x - 1. Therefore h(-20) = -20 - 1 = ... what?

OpenStudy (anonymous):

That's what I got wrong!! I thought it was: h(5)= 5-1 = 4 But, it's h(5)= -20-1 = -21!!

OpenStudy (anonymous):

h(x)**

OpenStudy (jamesj):

h(5) indeed is equal to 5 - 1 =4. But h(-20) = -20 - 1 = -21. Now you are trying to find the value of h(j(5)). That is NOT equal to h(5). First you have to evaluate j(5).

OpenStudy (jamesj):

In other words, \[ h(j(5)) \neq h(5) \] because \( j(5) \neq 5 \). Instead, as j(x) = -4x \[ j(5) = -4(5) = -20 \]

OpenStudy (jamesj):

Therefore, \[ h(j(5)) = h(-20) = -21 \]

OpenStudy (anonymous):

Thank you again... But, now I need help doing the fractions. If f(x) = x - 4 and g(x) = 1/2 x - 2, find f[g(-6)]. How do I do fractions in this?

OpenStudy (jamesj):

Is \[ g(x) = \frac{1}{2x - 2} \] ?

OpenStudy (anonymous):

No.... It's one 2 on the bottom and the "x" is next to the line.

OpenStudy (anonymous):

so, it's like (1/2) x-2....?

OpenStudy (jamesj):

\[ g(x) = (x/2) - 2 \] ?

OpenStudy (anonymous):

umm,.

OpenStudy (jamesj):

this is the same as (1/2)x - 2.

OpenStudy (anonymous):